i need help putting this into standard form.. 3(x-1)²=300
can you help?

Answer by ankor@dixie-net.com(22740)

3(x-1)²=300
:
3(x^2 - 2x + 1) = 300
:
3x^2 - 6x + 3 = 300
:
3x^2 - 6x + 3 - 300 = 0
:
y = 3x^2 - 6x - 297, the standard form
: